Circular No. 6242

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995af IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur (OCA), reports his discovery

of an apparent supernova (V = 20.5) on two films obtained on Sept. 1.07 and 2.08 UT by D. Albanese and himself with the OCA Schmidt telescope; the supernova is located at α = 23^{h}24^{m}53^{s}.45, δ = +41^{o}24′11′′.5 (equinox 2000.0), which is 5′′.4 west and 8′′.6 north of the spiral galaxy’s center (the galaxy itself of mag V ∼ 18.0). An OCA Schmidt film taken on Sept. 20.95 shows the new object at V = 21.7. SN 1995af was not visible on OCA Schmidt films taken in 1991 Aug. and Sept. (limiting mag B = 22.0) or on 1995 Aug. 5.10 (to B = 20.0), nor is anything visible at this location on the first and second Palomar Sky Survey photographs. A nearby star (V = 19) has position end figures 53^{s}.33, 05′′.6.

SUPERNOVA 1995ad IN NGC 2139 R. H. McNaught, Anglo-Australian Observatory, reports the follow-

ing precise position obtained on Sept. 29.8 UT: α = 6^{h}01^{m}06^{s}.13, δ = −23^{o}40′29′′.0 (equinox 2000.0); this is 24′′.5 west and 7′′.7 south of the galaxy’s center. Both SN 1995ad and a nearby foreground star were at (unfiltered) mag 15.6, the latter having position end figures 03^{s}.45, 39′53′′.4. Pollas reports a photograph taken with the OCA Schmidt telescope on Sept. 29.18 UT shows the supernova at mag ∼ 15 and provides the following position end figures for SN 1995ad: 06^{s}.17, 28′′.6. Pollas estimates McNaught’s foreground star (above) at mag V ∼ 17.5, and notes the presence of another nearby star (V ∼ 16.0) with position end figures 00^{m}55^{s}.11, 41′07′′.5.

SGR 1806–20 C. Kouveliotou, G. J. Fishman, C. A. Meegan, and K. Hagedon report

for the Compton Observatory BATSE Team: “We have detected transient soft-γ-ray emission from a region consistent with the direction of the ‘soft gamma repeater’ (SGR) 1806–20. The BATSE location is centered at α = 272^{o}, δ = −20^{o} and has an error radius of ∼ 4^{o}. The outburst was recorded on Sept. 30.72852 UT and consists of one pulse with a triangular shape and 60-ms duration (40 ms FWHM). The spectrum of the event has an upper energy cutoff of ∼ 100 keV, consistent with a typical SGR spectrum. Follow-up observations of the source region at other wavelengths are strongly encouraged. ”

1995 October 3 Daniel W. E. Green